A hospital keeps records of its emergency-room traffic. Those records indicate that, beginning at 6:00 P.M. on any given day, the elapsed time until the first patient arrives has an exponential distribution with parameter λ = 6.9, where time is measured in hours. Determine the probability that, beginning at 6:00 P.M. on any given day, the first patient arrives

a. between 6:15 P.M. and 6:30 P.M.

b. before 7:00 P.M.

c. given that the first patient doesn’t arrive by 6:15 P.M., determine the probability that she arrives by 6:45 P.M.


1
Expert's answer
2022-05-12T12:46:10-0400

Denote by "T" a random variable that has an exponential distribution. Its probability density function is: "f(x)=\lambda e^{-\lambda x}, x\geq0". "f(x)=0" for "x<0." "\lambda=6.9" is a parameter of the distribution. a. The aim is to find the probability: "P(\frac{1}{4}<T<\frac12)." We get: "P(\frac{1}{4}<T<\frac12)=\int_{\frac14}^{\frac12}6.9 e^{-6.9 x}dx=-e^{-6.9 x}|_{\frac14}^{\frac12}=e^{-6.9\cdot\frac14}-e^{-6.9\cdot\frac12}\approx0.1464".

b. The aim is to find "P(T<1)." We get: "P(T<1)=\int_{0}^{1}6.9 e^{-6.9 x}dx=-e^{-6.9x}|_0^1=1-e^{-6.9}\approx0.9990"

c. The aim is to find "P(T<\frac34|T>\frac14)". Using the definition of the conditional probability, we have: "P(T<\frac34|T>\frac14)=\frac{P(\frac14<T<\frac34)}{P(T>\frac14)}". Compute both probabilities in the latter formula: "P(\frac14<T<\frac34)=\int_{\frac14}^{\frac34}6.9e^{-6.9x}dx=-e^{-6.9x}|_{\frac14}^{\frac34}=e^{-6.9\cdot\frac14}-e^{-6.9\cdot\frac34}\approx0.1725".

"P(T>\frac14)=\int_{\frac14}^{+\infty}6.9e^{-6.9x}dx=-e^{-6.9x}|_{\frac14}^{+\infty}\approx0.1782."

We get: "\frac{P(\frac14<T<\frac34)}{P(T>\frac14)}=\frac{0.1725}{0.1782}\approx0.9680".

Answer: a. "P(\frac{1}{4}<T<\frac12)\approx0.1464", b. "P(T<1)\approx0.9990", c. "P(T<\frac34|T>\frac14)\approx0.9680". All values are rounded to "4" decimal places.